Use the given statistics to complete parts​ (a) and​ (b). Assume that the populations are normally distributed.

​(a) Test whether

mu 1μ1greater than>mu 2μ2

at the

alphaαequals=0.10

level of significance for the given sample data.​(b) Construct a

99%

confidence interval about

mu 1μ1minus−mu 2μ2.

Population 1

Population 2

n

28

20

x overbarx

46.6

42.4

s

4.7

10.2

Find the test statistic for this hypothesis test.

1.72

​(Round to two decimal places as​ needed.)

Determine the​ P-value for this hypothesis test.

?

​(Round to three decimal places as​ needed.)
Answer #1

SE = sqrt(4.7^2/28 + 10.2^2/20) = 2.4476

Test statistic,
t = (46.6 - 42.4)/2.4476
t = 1.72

df = 20 - 1 = 19
p-value = 0.051